OpenStudy (anonymous):

calculate the volume of a 0.206M HI solution that needed to reduce 22.5mL of a 0.374 KMnO4 solution ? help plzzzzz 10HI + 2KMnO4 + 3H2SO4 ------> 5i2 + 2MnSO4 = 8H2O

OpenStudy (abmon98):

is there any chemical equation listed in your question?

OpenStudy (anonymous):

Ooh sorry i forget it !! 10HI + 2KMnO4 + 3H2SO4 ------> 5i2 + 2MnSO4 = 8H2O

OpenStudy (abmon98):

its okay :) Calculate the number of moles=Concentration(mol/dm^3)*Volume(dm^3) We change ml to dm^3 by dividing by 1000 Number of moles KMnO4=0.374*22.5/1000=0.008415 10HI + 2KMnO4 + 3H2SO4 ------> 5i2 + 2MnSO4 = 8H2O 10 moles of HI are required to react with 2 moles of KMnO4 what if 0.008415 were used of KMnO4 how many moles are required by HI. Calculate the number of moles=Concentration(mol/dm^3)*Volume(dm^3) Rearrange your equation to find Volume Volume=Number of Moles/Concentrate the answer you will be getting is in dm^3 so make sure you multiply it by 1000

OpenStudy (abmon98):

M(HI)*V(HI)/M(KMnO4)*V(KMnO4)=HI/KMnO4 M(HI) represent the concentration of HI V(HI) represent the volume of HI M(KMnO4) represent the concentration of KMnO4 M(KMnO4) represent the volume of KMnO4 HI/KMnO4=Mole ratio which is 10/2 according to your chemical equation.

OpenStudy (anonymous):

thnxxx bro ^^

OpenStudy (abmon98):

your most welcome :)